OpenStudy (anonymous):

I'LL GIVE A MEDAL The length of the base of a triangle is 6.6 cm and its height is 4.9 cm. If the area of the triangle is half of the product of its base length and height, the area of the triangle is __________ cm2. (Round your answer to the appropriate significant digits)

Directrix (directrix):

A = (1/2) b*h A = (1/2)* (6.6)*(4.9) A = ? @ZoeMarae Post what you get and we can compare answers. Okay?